OpenStudy (anonymous):

In hw#10 problem #1, I am to find the product of exponentials based off of hw#9 problem #4. I am to use the striaght out configuration as reference. All the g(alpha)'s, in hw#9 prob #4, are based on a sitraight up configuration. Would straight out be my initial condition, g0? I would not need to change anything about the g(alphas) from hw#9 #4 because the alphas are generic, would I?

OpenStudy (anonymous):

The \[e ^{\xi \alpha}\] would be the \[g(\alpha)\] from hw#9 prob#4

OpenStudy (anonymous):

Something like that but I think in your notes he described how to find the "twist" associated with a transformation. You shouldn't just be taking the ln of your gs.

OpenStudy (anonymous):

\[\xi = \left(\begin{matrix}\hat g \omega \\ \omega\end{matrix}\right) \]

OpenStudy (anonymous):

It's different for different joints in general but for revolute joints it's. \[\left(\begin{matrix}q_i*\omega_i \\ \omega_i\end{matrix}\right)\]

OpenStudy (anonymous):

what is/does the twist do for/to the manipulator?

OpenStudy (anonymous):

Why is the twist different for different joints?

OpenStudy (anonymous):

What is a twist in general?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

No I'm asking you

OpenStudy (anonymous):

It is a way to discribe a turning at a joint.

OpenStudy (anonymous):

the twist does not explain a translation.

OpenStudy (anonymous):

Not correct. A twist is nothing more/nothing less than a velocity(linear and angular)\[\xi=\left(\begin{matrix}v_1 \\ v_2\\ v_3\\w_1\\w_2\\w_3\end{matrix}\right)\]For the 3D case

OpenStudy (anonymous):

When I think of velocity, I think of an object moving a distance in a period of time (car, baseball, person running ...). When I think of velocity in connection with athe manipulator arm I think of speed of rotation of the arm. This speed of rotation is always angular. I do not grasp the idea of the involvment of the linear velocity. Why do I have a v in the twist?

OpenStudy (anonymous):

It's not the velocity of the joint you're describing it's the velocity of the transformation from joint to joint. All the joints on your manipulator are revolute but clearly you can see that there is a linear velocity as well as an angular one of each link.

OpenStudy (anonymous):

Do you mean the speed at which any joint moves from that joints last position to its new position?

OpenStudy (anonymous):

I'm not following. A revolute joint by definition only revolves. However since it is attahed to a link of length greater than 0 that link both revolves and has a linear displaement(and therefore linear veloity). Just keep your elbow still and turn it. You will see that your hand both rotates and change position.

OpenStudy (anonymous):

I understand.

OpenStudy (anonymous):

The initial condition for #1 would be \[g _{0} = \left[\begin{matrix}R _{x}(-\pi/2) & \left(\begin{matrix}0 \\ l _{1}+l _{2}+l _{3}+l _{4} \\ l _{0}\end{matrix}\right) \\ 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

The individual twists would be base off of hw#9 prob#4

OpenStudy (anonymous):

What do you mean based off? I don't think there's any reason you have to look at HW9 for this problem.

OpenStudy (anonymous):

I thought the wording implied to use last weeks homework.

OpenStudy (anonymous):

Well you are doing the same thing as last week's homework but using a different method. So you should get the same results.

OpenStudy (anonymous):

I have class now. Thanks for your help. If your around I wiil be back at 12:30. Thanks.

OpenStudy (anonymous):

Ok that's fine.

OpenStudy (anonymous):

It looks like the joint between l3 and l4 has no rotation. Is that correct?

OpenStudy (anonymous):

It rotates about the y. The axis of rotation is coming out of the end effector.

OpenStudy (anonymous):

So my product of exponetials is \[g _{e}(\alpha) = e ^{\hat \xi _{1}\alpha _{1}}e ^{\hat \xi _{2}\alpha _{2}}e ^{\hat \xi _{3}\alpha _{3}}e ^{\hat \xi _{4}\alpha _{4}}e ^{\hat \xi _{5}\alpha _{5}}*g _{0} \]

OpenStudy (anonymous):

Yes I think that's right.

OpenStudy (anonymous):

I find the indivial values of \[\xi \] is what I find

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

q a point the represents the joint position

OpenStudy (anonymous):

q is a point on the axis of rotation

OpenStudy (anonymous):

Is there a difference? The axis of rotation runs through the joint and the joint is all I am interested in. Do I have the understanding correct?

OpenStudy (anonymous):

Ah yes there is a slight difference. You may be able to find an "easier" point on the axis of rotation than the joint itself. You also may not be able to. It depends.

OpenStudy (anonymous):

In the figure, say I am interested in joint/axis 2. The q for this joint is \[\sqrt{(l _{0})^{2}+(l _{1})^{2}}\]

OpenStudy (anonymous):

q is the straight line distance from the origin to the joint.

OpenStudy (anonymous):

No that's not right...that's not the axis of rotation.

OpenStudy (anonymous):

The \[\omega \] at alpha2 is \[\left(\begin{matrix}0 \\ l _{1} \\ l _{0}\end{matrix}\right) \]

OpenStudy (anonymous):

No that's also not right. w is a unit vector aligned with the axis of rotation in the base frame.

OpenStudy (anonymous):

The q at alpha2 is \[\left(\begin{matrix}0 \\ l _{1} \\ l _{0}\end{matrix}\right) \]

OpenStudy (anonymous):

Yes that's the easiest one.

OpenStudy (anonymous):

I take q and turn it into q-hat

OpenStudy (anonymous):

q is a 3x1, q-hat is a 3x3 matrix

OpenStudy (anonymous):

How do I come up with omega

OpenStudy (anonymous):

how are you going to make a 3x3 matrix out of a 3x1 point?

OpenStudy (anonymous):

\[q = \left(\begin{matrix}a \\ b \\ c\end{matrix}\right) \] \[q-\hat = \left[\begin{matrix}0 & c & -b\\ -c & 0 & a \\ b & -a & 0\end{matrix}\right] \]

OpenStudy (anonymous):

I don't think that's quite right. What you're actually doing is qxw. Here is how to make a matrix out of q if that's how you want to compute it: http://en.wikipedia.org/wiki/Cross_product#Conversion_to_matrix_multiplication w is just a unit vector pointing in the same direction as the axis of rotation.

OpenStudy (anonymous):

So if my axis of rotation is y then w = (0 1 0) transpose

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

I thought anything with a hat on it was suppose to be some form of even matrix. I do not want to make this harder on myself by doing something that we were not shown. If q is a 3x1 and w is a 3x1 and I do remember to do the cross product to get qw then I am fine.

OpenStudy (anonymous):

Yeah that's fine. Things with a hat are matrices yes. There is a technique for finding a cross product which may be easier than whatever you know which is to make that matrix and then just multiply it by the w. It's your choice but it's really just a cross product.

OpenStudy (anonymous):

For #2a, I would use the Pieper's Solution.

OpenStudy (anonymous):

I think Pieper's solution is just the one that breaks down the inverse kinematics into the hand and the wrist. You can just use geometry for 2

OpenStudy (anonymous):

For #2a, I find the angles and create a forward kinematics equation.

OpenStudy (anonymous):

You don't need to find the angles. You just need to compute the forward kinematics equation.

OpenStudy (anonymous):

\[g _{e}(\alpha) = g _{1}(\alpha _{1})g _{2}(\alpha _{2})\]

OpenStudy (anonymous):

Ok sure but write it out interms of ls and cosines and sines.

OpenStudy (anonymous):

This is what I have \[g _{e}(\alpha) = \left[\begin{matrix}\cos (\alpha _{1}) & -\sin (\alpha _{1}) & 0\\ \sin (\alpha _{1}) & \cos (\alpha _{1}) & 0 \\ 0 & 0 & 1\end{matrix}\right]*\left[\begin{matrix}\cos (\alpha _{2}) & -\sin (\alpha _{2}) & l _{1}\\ \sin (\alpha _{2}) & \cos (\alpha _{2}) & 0 \\ 0 & 0 & 1\end{matrix}\right]*\left[\begin{matrix}1 & 0 & l _{2}\\ 0 & 1 & 0 \\ 0 & 0 & 1\end{matrix}\right] \]

OpenStudy (anonymous):

I would just write it out in equations. It should look familiar. It's been done a couple times in the past.

OpenStudy (anonymous):

I think it was done in HW1 if you look back. In this case the intuitive solution is easier than this product of Lie algebras.

OpenStudy (anonymous):

So when it says planar, that is another way of saying 2D forward kinematics?

OpenStudy (anonymous):

If that is the case, then when asked for \[d _{e} = \left(\begin{matrix}x _{e} \\ y _{e}\end{matrix}\right)\] xe would be an equation with l's and cos, and ye would have l's and sin

OpenStudy (anonymous):

That would mean \[g _{e}(\alpha) = d _{e}(\alpha)\]

OpenStudy (anonymous):

in 2D

OpenStudy (anonymous):

I'm not following you here. Yes planar means enclosed in a plane or equivalently 2D. Yes xe is ls and cosines and ye is ls and sines. I don't know what you mean by g=d

OpenStudy (anonymous):

I meant to write qe not ge. qe = de

OpenStudy (anonymous):

For #2b, I just use the geometric approach to find alpha1 and alpha2. For #2c, I look at the equations that solve for alpha1 and alpha2 and find the cases where no solution exists. Am I correct in the approach in solving both questions?

OpenStudy (anonymous):

If your approach makes sense to you then you should feel confident in it.

OpenStudy (anonymous):

Can you explain #3? I think I am looking for the inverse kinematics but I will be using the geometric method to find the the position and the Pieper's method to find the angles. How is this manipulator fully controllable as opposed to all the other manipulators we have dealt with? What does fully controllable have to do with using the Pieper's method?

OpenStudy (anonymous):

I'm actually not really sure what is meant by fully controllable and how that relates to pieper's method. What difficulty are you having in understanding the rest of the problem?

OpenStudy (anonymous):

#3 starts out asking to 'Work out' the inverse kinematics and tells me to find the solution for the plane using the Pieper's method and find the position part using the geometric method. The second paragraph talks about solving the problem by spliting the forward kinematics into the gw and gh parts. Giving gw = (ge*)(g4)^-1. This is as much as I understand the question to be asking.

OpenStudy (anonymous):

Yes that is what the question is asking. The main goal is to solve the inverse kinematics using Pieper's method. Your solution from problem 2 should be of use. That's all he's saying.

OpenStudy (anonymous):

In #2, I solved the forward kinematics with two equations. xe = l's and cos and ye = l's and sin.

OpenStudy (anonymous):

In 2b you solved inverse kinematics

OpenStudy (anonymous):

Solving for inverse kinematics is just finding the alphas.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Can I not do that in #3 without going through the gw and gh steps?

OpenStudy (anonymous):

You can try but I think you will find it's hard.

OpenStudy (anonymous):

\[g _{w} = g _{e}^{*}g _{4}^{-1} = g _{1}(\alpha _{1})g _{2}(\alpha _{2})g _{3}\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

What does \[g _{e}^{*}\] mean

OpenStudy (anonymous):

Desired end effector configuration.

OpenStudy (anonymous):

So I am suppose to have some known end effector configuration? I do not have a known end effector configuration unless I work out the forward kinematics. Would I use the form I used in #2a (l's, cos and sin) or would I use the homogen. form?

OpenStudy (anonymous):

State the problem of inverse kinematics.

OpenStudy (anonymous):

I do not follow.

OpenStudy (anonymous):

When someone asks you to work out inverse kinematics, what are they asking?

OpenStudy (anonymous):

They want me to find the angles of the manipulator joints.

OpenStudy (anonymous):

For a given end effector position.

OpenStudy (anonymous):

In problem 2a, you solved forward kinematics(end effector configuration for given alpha) even though you were not actually given the alphas. You solved it symbolically such that any alphas could be plugged in. This is exactly the same thing, except what is given and what is desired is reversed.

OpenStudy (anonymous):

I follow what you are saying. I am having difficulty doing it.

OpenStudy (anonymous):

You wrote the equation for gw right? Using this equation and the given information you can figure out the position of the wrist.

OpenStudy (anonymous):

I start with \[g _{e}^{*}g _{4}^{-1} = g _{1}(\alpha _{1})g _{2}(\alpha _{2})g _{3}\] I know \[g _{e}(\alpha) = g _{1}(\alpha _{1})g _{2}(\alpha _{2})g _{3}(\alpha _{3})g _{4}\]

OpenStudy (anonymous):

I don't think you understood what I was saying. ge* is given. g4 is given. You can get gw.

OpenStudy (anonymous):

g4 is a translation over the distance l3.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

If I multiply a translation by the inverse of that translation would I get a unit vector?

OpenStudy (anonymous):

as in g*g^-1 = g1(alpha1)g2(alpha2)g3(alpha3)g4 * (g4)^-1

OpenStudy (anonymous):

Yes of course multiplying a matrix by its inverse gives you the identity. What do you mean?